OpenStudy (anonymous):

Find the ratio of x to y x/5=2/3=5/y 2/3 4/9 1

OpenStudy (anonymous):

x/5 = 2/3 => x = 10/3 5/y = 2/3 => y = 15/2. =====> x/y = 4/9